find the equation of the parallel line to 6x + 3y=15 that passes through points (6,9. write in slope-intercept form.
Answer by palanisamy(496)

The equation of the given line to 6x + 3y= 15
Let the equation of the parallel line be 6x + 3y = c ...(1)
This line passe through the point (6,9)
Therefore, 6*6+3*9 = c
] c = 63
So The equation of the required line is 6x+3y = 63
3y = -6x+63
y = -2x+21